OpenStudy (anonymous):

Please help!!!!! survey was conducted among 200 students of age groups 7–12 years and 13–18 years to find their favorite movie genre. The students had to select any one genre out of comedy, action, and adventure. Out of the 100 students in the age group 7–12 years who participated in the survey, 55 liked action or adventure movies. The total number of students of both age groups who liked comedy was 98. Using a two-way table, compute the total number of students in the age group 13–18 years who liked action or adventure movies.

OpenStudy (anonymous):

Please explain?!

OpenStudy (victoriasushchik):

ok so The total number of students of both age groups who liked comedy was 98. so 98/2= 49 And it says already that the 7-12 age group has 55 liked action or adventure. And 55+49=104 But there's 100 people for each age group so 104-100=4 and 49+4=53 answer:53

OpenStudy (anonymous):

Ohh okay, thank you so much!!

OpenStudy (victoriasushchik):

np =)